A galvanic cell is powered by the following redox reaction: 


O₂(g)+4 H⁺(a q)+2 Zn(s) → 2H₂O(i)+2Zn²⁺(aq)


Answer the following questions about this cell. If you need any electrochemical data, be sure you get it from the ALEKS Data tab. 

  • Write a balanced equation for the half-reaction that takes place at the cathode. 

  • Write a balanced equation for the half-reaction that takes place at the anode. 

  • Calculate the cell voltage under standard conditions. 
Answer #1

A: Anode: Increase in oxidation no. = Oxidation
C: Cathode : Decrease in oxidation no. = Reduction

Anode is an ELECTRODE where OXIDATION occurs. (loss of electrons)
HALF REACTION THAT TAKES PLACE AT ANODE:-

Zn(s) → Zn2+(aq) + 2e-

HALF REACTION THAT TAKES PLACE AT CATHODE :-

O2(g) + 4H+(aq) + 4e– → 2H2O(l)

CELL VOLTAGE UNDER STANDARD CONDITIONS :-

O2(g) + 4H+(aq) + 4e– → 2H2O(l) E0 = 1.23 V (this is the most likely reduction (largest positive E0) . So other reaction will be reversed)
Zn(s) → Zn2+(aq) + 2e– - E0 = + 0.76V
[Zn(s) → Zn2+(aq) + 2e–] x 2 - E0 = + 0.76 V

The equations are then added
O2(g) + 4H+(aq) + 4e- + 2Zn(s) → 2H2O(l) + 2Zn2+(aq) + 4e-

O2(g) + 4H+(aq) + 2Zn(s) → 2H2O (l) + 2Zn2+(aq)

emf = 1.23 V + (+ 0.76 V) = 1.99 V